# Vec2Face: Scaling Face Dataset Generation with Loosely Constrained Vectors

### v1

@@ Abstract
This paper studies how to synthesize face images of non-existent persons, to create a dataset that allows effective training of face recognition (FR) models. Two important goals are (1) the ability to generate a large number of distinct identities (inter-class separation) with (2) a wide variation in appearance of each identity (intra-class variation). However, existing works 1) are typically limited in how many well-separated identities can be generated and 2) either neglect or use a separate editing model for attribute augmentation. We propose Vec2Face, a holistic model that uses only a sampled vector as input and can flexibly generate and control face images and their attributes. Composed of a feature masked autoencoder and a decoder, Vec2Face is supervised by face image reconstruction and can be conveniently used in inference. Using vectors with low similarity among themselves as inputs, Vec2Face generates well-separated identities. Randomly perturbing an input identity vector within a small range allows Vec2Face to generate faces of the same identity with robust variation in face attributes. It is also possible to generate images with designated attributes by adjusting vector values with a gradient descent method. Vec2Face has efficiently synthesized as many as 300K identities with 15 million total images, whereas 60K is the largest number of identities created in the previous works. FR models trained with the generated HSFace datasets, from 10k to 300k identities, achieve state-of-the-art accuracy, from 92% to 93.52%, on five real-world test sets. For the first time, our model created using a synthetic training set achieves higher accuracy than the model created using a same-scale training set of real face images (on the CALFW test set). Code link: https://haiyuwu.github.io/vec2face.github.io/
@@ 1 Introduction
We aim to synthesize face images in a way that enables large-scale training sets for FR models, which have the potential to address privacy issues arising with web-scraped datasets of real face images. It is generally recognized that a good training set for FR should have high inter-class separability (Kim et al., 2023; Boutros et al., 2023a) and intra-class variation (Qiu et al., 2021; Boutros et al., 2022a; Kim et al., 2023; Boutros et al., 2023a).
@@
However, existing methods lack flexibility in controlling the generation process, leading to unsatisfactory inter-class and intra-class results. On the one hand, Boutros et al. (2023a) point out that identities generated in (Qiu et al., 2021; Boutros et al., 2022b; a; 2024) have relatively low separability because identity generation is controlled by either a single coefficient or by hard class labels. On the other hand, while it is useful to employ identity features as a condition to increase inter-class separability (Boutros et al., 2023a; Papantoniou et al., 2024), it is non-trivial to increase intra-class variation without the use of a separate model such as ControlNet (Zhang et al., 2023) or a model for style transfer.
@@
In light of the above, this paper proposes Vec2Face, a flexible model enabling the generation of large-scale FR datasets. In inference, the workflow of Vec2Face is similar to (Papantoniou et al., 2024; Boutros et al., 2023a), which use a random vector as input and output a face image. But the interesting property of Vec2Face is that a small perturbation of the input vector leads to a face image of the same identity but with small changes in appearance, e.g., pose, expression, and facial hair, while a large perturbation of the input vector produces a face image of a different identity (Fig. 1). With this property, and because of the high-dimensionality (i.e., 512-dim) of the input vectors, we can easily 1) sample a large number of identity vectors by controlling their similarity with each other to be under a certain threshold and 2) synthesize a large number of face images by perturbing the identity vectors within a small range. Additionally, we can guide the image content at time of generation using a gradient descent method to efficiently generate faces with minority attributes, such as extreme pose angles.
@@
Using a proper cosine similarity threshold (i.e., 0.3) to sample vectors, the special design of Vec2Face ensures a high ratio of transferring the similarity between vectors to the similarity between generated images, resulting in large inter-class separability; intra-class variation is also properly large because we controlled the variance of the noise for perturbation, so the generated images have $\geq$0.5 similarity with the identity vector, exhibiting various appearances while preserving the identity. Experimentally, we verify these inter-class and intra-class variations are beneficial (in hyper-parameter analysis). Because of this, FR models trained on the synthesized 10k identities and 500k images demonstrate very competitive accuracy compared with the state-of-the-art synthetic databases of the same scale (Kim et al., 2023; Papantoniou et al., 2024). Importantly, we show that scaling up the synthetic training data to 10 million images and 200k identities leads to further accuracy improvement. It is also worth noting that results on CALFW (Zheng et al., 2017) represent the first instance where a model trained with synthetic data yields higher accuracy than that trained with real data at the same scale. To our knowledge, this is the first time a synthetic dataset can be this large and more useful.
@@
The training of Vec2Face aims to let a vector generate a face and let its perturbation magnitudes reflect face change magnitudes. To this end, in training we use vectors extracted from real face images by a pre-trained FR model. These vectors are fed into a feature masked autoencoder architecture, after which a decoder produces images similar to the real ones on the pixel level. We summarize the main points of this paper below.
@@
  • We propose Vec2Face, a face synthesis model that allows us to efficiently synthesize a large number of face images and identities. Under an appropriate similarity threshold, Vec2Face effectively increases inter-class separability and intra-class variations of synthesized face images.
@@
  • With a generated dataset of 10k identities and 500k images, we achieve state-of-the-art FR accuracy measured as the average on five real-world test sets compared with existing synthetic datasets. Our performance even surpasses the accuracy obtained by real training data on one of the five test sets. We also report that scaling to larger synthetic training data leads to further improvement.
@@ 2 Related Work
Figure 1: Sample images generated by Vec2Face. From a random vector we generate a face (ID image). We then add random values to this vector to generate diverse face images. Larger perturbation added to this vector results in larger dissimilarity to the ID images. The images in the frame are likely the same people as the ID image.
@@
Class conditioned image generation. VQGAN (Esser et al., 2021), MAGE (Li et al., 2023), DiT (Peebles & Xie, 2023), U-ViT (Bao et al., 2022), MDT (Gao et al., 2023), and VAR (Tian et al., 2024) control the object classes of generated images using discrete class labels as a condition, and so do not support the generation of new classes. Therefore, they are not applicable to scalable face/identity generation. Different from them, Vec2Face uses continuous face features and thus technically can generate an unlimited number of new identities given appropriate feature sampling.
Feature conditioned image generation. The models trained with conditioned features, either from a text encoder or an image encoder, perform impressively on diverse class generation. However, (Papantoniou et al., 2024) reports that text embedding exhibits a deficiency in face identity retention. (Chen et al., 2023; Li et al., 2024; Xiao et al., 2023; Wang et al., 2024; Valevski et al., 2023; Ye et al., 2023; Yan et al., 2023; Papantoniou et al., 2024) use an image encoder, either CLIP (Radford et al., 2021) or a FR model, to guide the generation process. These methods are less efficient and effective because 1) the image generation is computationally expensive on these models and 2) they use a fixed vector to generate the images from an identity, which loses the control over the attribute variation. We posit that face feature vectors contain both identity and attribute information. To this end, our method learns image generation and achieves output control purely on face features.
Synthetic face image datasets. Existing approaches are primarily based on two methods: GAN-based models and diffusion models, with the latter generally producing datasets that enable FR models to achieve better accuracy. Here, we briefly summarize some important works.
@@
GAN-based methods: SynFace (Qiu et al., 2021), Usynthface (Boutros et al., 2022b), SFace (Boutros et al., 2022a), SFace2 (Boutros et al., 2024), and ExfaceGAN (Boutros et al., 2023b) leverage pre-trained GAN models to generate datasets. These methods are less effective because the pre-trained GAN models were not trained in an identity-aware manner. Departing from them, our method explicitly uses face features in training, so the generated images are identity-aware.
@@
Diffusion model-based methods: DCFace (Kim et al., 2023) uses a strong pre-trained diffusion model (Choi et al., 2021) and an additional style-transfer model to increase the identity separability and attribute variation. Again, this pretrained model is also not identity-aware. IDiff-Face (Boutros et al., 2023a) combines the pretrained encoder and decoder from VQGAN (Esser et al., 2021) with a latent diffusion model conditioned by identity features to control the separability of the generated identities. Although the latent diffusion model is identity-aware, the pre-trained VQGAN reduces the sensitivity of identity on the output. Arc2Face (Papantoniou et al., 2024) fine-tunes a pre-trained stable diffusion model (Rombach et al., 2022) on the WebFace42M dataset to increase the generalizability on face. It combines the identity features and CLIP (Radford et al., 2021) to control the identity of output images. The pose variation is increased by adding an additional ControlNet (Zhang et al., 2023). Neither the pre-trained stable diffusion model nor CLIP are optimized for face and the slow processing speed strongly limits large scale dataset generation. On the contrary, our method is specifically designed for face dataset generation and is both identity-aware and fast in generation. This allows our method to achieve state-of-the-art performance while easily scaling the dataset size to 15M.
@@ 3 Method
3.1 Vec2Face
As shown in Fig. 2, Vec2Face consists of a FR model, a feature Masked Autoencoder, an image decoder, and a patch-based discriminator (Isola et al., 2017; Yu et al., 2022).
@@ 3.1 Vec2Face
Feature masked auto-encoder (fMAE). Similar to MAE (He et al., 2022), the model is forced to learn better information by masking out the input. Different from the MAE that masks an input image and introduce mask tokens to form the full-size image, the fMAE uses a 2D feature map as input and masks this feature map. To provide more useful information, the image feature is used to form the full-size feature map. Specifically, the rows in the feature map are randomly masked out by $x\%$ before the encoding process, where $x\%\in\mathcal{N}_{truncated}(max=1,min=0.5,mean=0.75)$, and the image feature is filled in the masked out positions to form the full-size feature map before being processed by the decoder. The structure is in the Appendix.
Image Decoder. Finally, the new feature map is passed through a simple image decoder to generate/reconstruct the images. To improve image quality, a patch-based discriminator is integrated to form a GAN-type training.
Figure 2: Architecture of Vec2Face. Given a real face image, we compute its feature “IM feature” using a face recognition (FR) model. The feature is expanded into a feature map, and the latter is processed by a feature masked auto-encoder (fMAE). Inside the fMAE, the rows in the feature map are randomly masked out before being processed by the encoder. The image features are then introduced to form the full-size feature map before being processed by the decoder. Finally, the decoder outputs a feature map. Outside the fMAE, a small image decoder reconstructs the pixels in the original image based on the output of fMAE. During inference, Vec2Face accepts a randomized vector and generates a face image. This process has properties demonstrated in Fig. 1.
Loss function. The training objective functions include an image reconstruction loss and an identity loss. The reconstruction loss can be written as,
@@
which compares the pixel-level difference between the reconstructed image $IM_{rec}$ and the input ground truth images $IM_{gt}$. The identity loss is written as:
@@
meaning that the features of the reconstructed image and original image extracted using the FR model should be close. In addition, the perceptual loss (Zhang et al., 2018) is used to ensure the correct face structure at the early training stage and a patch-based discriminator (Isola et al., 2017; Yu et al., 2022) is used to form the GAN loss to increase the sharpness of the generated images. The training loss is:
@@
Without $L_{id}$, Vec2Face has slightly degraded generation performance but still works reasonably well on identity preservation. Without $L_{lpips}$, the reconstructed images do not have clear face structures. Using $L_{GAN}$ at early training stages makes the model convergence difficult, so it is used only after 1,000 epochs. Examples of loss effect on image reconstruction are in Appendix.
Sample generation results. Image generation results are shown in Fig. 3. For both seen and unseen identities, the generated images have very similar identities to the original images. Moreover, the reconstructed image would remove image border artifacts and some complex backgrounds and translate sketches into photo-realistic images.
Figure 3: Generation results for training (left) and unseen (right) identities. Images in the same row are of the same identity. We extract features from original images and feed them to Vec2Face. We observe that the reconstructed images still maintain the same identity while removing some image borders and backgrounds and transferring sketches into photo-realistic images.
3.2 Sampling identities and face images
For Vec2Face, dissimilar input vectors will likely lead to images of different identities, allowing us to generate different identity images; controlling the perturbed identity vectors in a proper similarity range with the identity vector can generate face images with different attributes while preserving the identity.
Sampling identity images. Suppose we want to generate $n$ identity images from $n$ identity vectors $v_{1},v_{2},...,v_{n}$, where $v_{i}\in\mathbb{R}^{d},i=1,2,...,n$. We ask that the similarity between any pair of identity vectors be less than a threshold $\tau$: $sim(v_{i},v_{j})\leq\tau,\hskip 5.69054pt\forall\;i\neq j$. This ensures that the generated images are of different identities. This paper follows (Papantoniou et al., 2024) in using PCA, without dimension reduction, learned on identity features of MS1MV2 (Deng et al., 2019) to randomly sample vectors. Specifically, we first compute the mean and covariance of the PCA-transformed identity features, then use these statistics to generate new samples in the PCA space via multivariate normal distribution, and finally inverse transform these samples back to the original feature space to obtain new identity vectors. Newly sampled vectors that have similarity $\leq\tau=0.3$ with all existing vectors are kept, otherwise they are dropped. Interestingly, we notice that, in the high-dimensional (512 in this paper) feature space, most of the randomly sampled vectors fit the condition due to sparsity. Thus, in practice, only 1.7% of the sampled 300k vectors are filtered out.
Sampling images for each identity. After obtaining $n$ identity vectors $v_{1},v_{2},...,v_{n}$, we sample $m$ vectors for each identity vector by perturbation:

### v4

@@ Abstract
This paper studies how to synthesize face images of non-existent persons, to create a dataset that allows effective training of face recognition (FR) models. Besides generating realistic face images, two other important goals are: 1) the ability to generate a large number of distinct identities (inter-class separation), and 2) a proper variation in appearance of the images for each identity (intra-class variation). However, existing works 1) are typically limited in how many well-separated identities can be generated and 2) either neglect or use an external model for attribute augmentation. We propose Vec2Face, a holistic model that uses only a sampled vector as input and can flexibly generate and control the identity of face images and their attributes. Composed of a feature masked autoencoder and an image decoder, Vec2Face is supervised by face image reconstruction and can be conveniently used in inference. Using vectors with low similarity among themselves as inputs, Vec2Face generates well-separated identities. Randomly perturbing an input identity vector within a small range allows Vec2Face to generate faces of the same identity with proper variation in face attributes. It is also possible to generate images with designated attributes by adjusting vector values with a gradient descent method. Vec2Face has efficiently synthesized as many as 300K identities, whereas 60K is the largest number of identities created in the previous works. As for performance, FR models trained with the generated HSFace datasets, from 10k to 300k identities, achieve state-of-the-art accuracy, from 92% to 93.52%, on five real-world test sets (i.e., LFW, CFP-FP, AgeDB-30, CALFW, and CPLFW). For the first time, the FR model trained using our synthetic training set achieves higher accuracy than that trained using a same-scale training set of real face images on the CALFW, IJBB, and IJBC test sets.
@@ 1 Introduction
We aim to synthesize face images in a way that enables large-scale training sets for FR models, which have the potential to address privacy issues arising with web-scraped datasets of real face images (DeAndres-Tame et al., 2024; Melzi et al., 2024; Shahreza et al., 2024). It is generally recognized that a good training set for FR should have high inter-class separability (Kim et al., 2023; Boutros et al., 2023a) and proper intra-class variation (Qiu et al., 2021; Boutros et al., 2022a; Kim et al., 2023; Boutros et al., 2023a).
@@
However, existing methods lack flexibility in controlling the generation process, leading to unsatisfactory inter-class and intra-class results. On the one hand, (Boutros et al., 2023a) points out that identities generated in (Qiu et al., 2021; Boutros et al., 2022b; a; 2024) have relatively low separability because identity generation is controlled by either a single coefficient (Deng et al., 2020) or by hard class labels. On the other hand, while it is useful to employ identity features as a condition to increase inter-class separability (Boutros et al., 2023a; Papantoniou et al., 2024), it is non-trivial to increase intra-class variation without the use of a separate model such as ControlNet (Zhang et al., 2023) or a model for style transfer.
@@
In light of the above, this paper proposes Vec2Face, a holistic model enabling the generation of large-scale FR datasets. In inference, the workflow of Vec2Face is similar to (Papantoniou et al., 2024; Boutros et al., 2023a), which uses a random vector as input and outputs a face image. But the interesting property of Vec2Face is that a small perturbation of the input vector leads to a face image of the same identity but with small changes in appearance, e.g., pose, age, and facial hair and that a large perturbation of the input vector produces a face image of a different identity (Fig. 1). With this property, and because of the high-dimensionality (i.e., 512-dim) of the input vectors, we can easily 1) sample a large number of well-separated identity vectors by controlling their similarity with each other to be under a certain threshold and 2) synthesize a large number of face images with various attributes by perturbing the identity vectors within a small range. Additionally, we can guide the image content at time of generation using a gradient descent method to efficiently generate faces with designated attributes, such as extreme pose angles and good image quality.
@@
Using a proper cosine similarity threshold (i.e., 0.3) to constrain the identity vector sampling, Vec2Face transfers the similarity between vectors to the similarity between generated images, resulting in large inter-class separability; intra-class variation is also properly large because we control the variance of the noise for perturbation, so the perturbed identity vectors have $\geq$0.5 similarity with the identity vector, exhibiting various appearances while preserving the identity on the generated images. Experimentally, we verify these inter-class and intra-class variations are beneficial to the FR model performance (in Section 4.4). Because of this, FR models trained on the synthesized 10K identities and 500K images demonstrate very competitive accuracy compared with the state-of-the-art synthetic databases of the same scale (Kim et al., 2023; Papantoniou et al., 2024). Importantly, we show that scaling up the synthetic training data to 15 million images and 300K identities leads to further accuracy improvement. It is also worth noting that results on CALFW (Zheng et al., 2017), IJBB (Whitelam et al., 2017), and IJBC (Maze et al., 2018) represent the first instance where a model trained with synthetic data yields higher accuracy than that trained with real data at the same scale. To our knowledge, this is the first time a synthetic dataset can be this large and useful.
@@
The training of Vec2Face aims to let a vector generate a face and let its perturbation magnitudes reflect face change magnitudes. To this end, in training we use vectors extracted from real face images by a pre-trained FR model. These vectors are fed into a feature masked autoencoder architecture, after which an image decoder produces images similar to the real ones at the pixel level. We summarize the main points of this paper below.
@@
  • We propose Vec2Face, a face synthesis model that allows us to efficiently synthesize a large number of face images and well-separated identities. Under an appropriate similarity threshold, Vec2Face effectively increases inter-class separability and intra-class variation of synthesized face images.
@@
  • With a generated dataset of 10K identities and 500K images, we achieve state-of-the-art FR accuracy measured as the average on five real-world test sets compared with existing synthetic datasets. Our performance even surpasses the accuracy obtained by real training data on three test sets. We also report that scaling to larger synthetic training data leads to further improvement.
@@ 2 Related Work
Figure 1: Example images generated by Vec2Face. From a random vector, we generate a face (ID image). We then perturb this vector with random values to generate diverse face images. Larger perturbation added to this vector results in larger dissimilarity to the ID images. The images in the frame are likely of the same people as the ID image.
@@
Discrete class label conditioned image generation. VQGAN (Esser et al., 2021), MAGE (Li et al., 2023), DiT (Peebles & Xie, 2023), U-ViT (Bao et al., 2022), MDT (Gao et al., 2023), and VAR (Tian et al., 2024) control the object classes of generated images using discrete class labels as condition. Therefore, these methods do not support the new class generation and are not applicable to the context of scalable face identity generation. Differently, Vec2Face uses continuous face features as input and thus technically can generate infinite new identities given appropriate feature sampling.
Identity feature conditioned image generation. Using a static feature as condition to guide the model to generate images for one identity is another approach. (Xiao et al., 2023; Chen et al., 2023; Li et al., 2024b) use CLIP (Radford et al., 2021) to encode the face embedding, but this representation is limited by CLIP’s ability on facial encoding (Papantoniou et al., 2024). To improve the identity fidelity of the generated images, (Wang et al., 2024; Valevski et al., 2023; Ye et al., 2023; Wood et al., 2021; Papantoniou et al., 2024) use identity features extracted by a FR model as condition. However, an external model (e.g., ControlNet (Zhang et al., 2023)) is needed to increase the attribute variation. Different from them, Vec2Face generates images for one identity using dynamic perturbed identity features that have high similarity value with their identity feature. This keeps a high identity fidelity while providing a way to control the attribute variation without using any external models.
@@
Synthetic face image datasets. Existing approaches are primarily either GAN-based or diffusion-based. For the former, SynFace (Qiu et al., 2021), Usynthface (Boutros et al., 2022b), SFace (Boutros et al., 2022a), SFace2 (Boutros et al., 2024), and ExfaceGAN (Boutros et al., 2023b) leverage pre-trained GAN models to generate datasets. These methods are less effective because the pre-trained GAN models were not trained in an identity-aware manner. Departing from them, our GAN-based method explicitly uses face features in training, so the generated images are identity-aware.
@@
For diffusion model-based methods, DCFace (Kim et al., 2023) uses a strong pre-trained diffusion model (Choi et al., 2021) and an additional style-transfer model to increase the identity separability and attribute variation. Again, this pre-trained model is not identity-aware. IDiff-Face (Boutros et al., 2023a) combines the pre-trained encoder and decoder from VQGAN (Esser et al., 2021) with a latent diffusion model conditioned by identity features to control the separability of the generated identities. Although the latent diffusion model is identity-aware, the pre-trained VQGAN compromises such effect. Arc2Face (Papantoniou et al., 2024) fine-tunes a pre-trained stable diffusion model (Rombach et al., 2022) on the WebFace42M dataset to increase the generalizability on face. It combines the FR feature and CLIP (Radford et al., 2021) feature to control the identity of output images. The pose variation is increased by adding an additional ControlNet (Zhang et al., 2023). However, neither the pre-trained stable diffusion model nor CLIP are optimized for face, and the slow processing speed strongly limits large scale dataset generation. Conversely, our method is specifically designed for face dataset generation and is efficient and identity-aware. This allows our method to easily scale the dataset size to 15 million images and achieve state-of-the-art FR accuracy.
@@ 3 Method
3.1 Vec2Face: Architecture and Loss Function
As shown in Fig. 2, Vec2Face consists of a pretrained FR model, a feature masked autoencoder (fMAE), an image decoder, and a patch-based discriminator (Isola et al., 2017; Yu et al., 2022).
@@ 3.1 Vec2Face: Architecture and Loss Function
Feature masked auto-encoder (fMAE). Similar to MAE (He et al., 2022), the model is forced to learn better representations by masking out the input. Different from the MAE that masks an input image and introduces mask tokens to form the full-size image, the fMAE uses a 2D feature map as input and masks this feature map. To provide more useful information, the projected image feature is used to form the full-size feature map. Specifically, the rows in the feature map are randomly masked out by $x\%$ before the encoding process, where $x\%\in\mathcal{N}_{truncated}(max=1,min=0.5,mean=0.75)$, and the projected image feature is filled in the masked out positions to form the full-size feature map before being processed by the decoder. The structure is in the Appendix A.9.
Image decoder and patch-based discriminator. Finally, the new feature map is passed through a simple image decoder to generate/reconstruct the image. To improve image quality, a patch-based discriminator (Isola et al., 2017; Yu et al., 2022) is integrated to form a GAN-type training.
Figure 2: Architecture of Vec2Face. Given a real face image, we compute its feature “IM feature” using a face recognition model. This feature is projected and expanded into a feature map, and the latter is processed by a feature masked autoencoder (fMAE). Inside the fMAE, the rows in the feature map are randomly masked out before being processed by the encoder. The projected image feature is then used to form the full-size feature map before being processed by the decoder. Finally, the decoder outputs a feature map. Outside the fMAE, a small image decoder reconstructs the pixels in the original image based on the output of fMAE. During inference, Vec2Face accepts a randomized vector and generates a face image. This process has properties demonstrated in Fig. 1.
Loss function. The training objective function includes an image reconstruction loss, an identity loss, a perceptual loss and a GAN loss. The reconstruction loss can be written as,
@@
which compares the pixel-level difference between the reconstructed image $IM_{rec}$ and the ground truth images $IM_{gt}$. The identity loss is written as:
@@
meaning that the features of the reconstructed image and original image extracted using the FR model should be close. In addition, the perceptual loss (Zhang et al., 2018) is used to ensure the correct face structure at the early training stage and a patch-based discriminator (Isola et al., 2017; Yu et al., 2022) is used to form the GAN loss to increase the sharpness of the generated images. The total loss is:
@@
Some findings of the loss function’s effects are: 1) without $L_{id}$, Vec2Face has slightly degraded generation performance but still works reasonably well on identity preservation; 2) without $L_{lpips}$, the reconstructed images do not have clear face structures; 3) using $L_{GAN}$ at early training stage makes the model convergence difficult, so it is used only after 1,000 epochs (80 is the choice after the code optimization). Examples of loss effects on image reconstruction are in Appendix A.2.
Image reconstruction results. Fig. 3 shows that, for both seen and unseen identities, the generated images have very similar identities to the original images. Moreover, the reconstructed image would remove image border artifacts and some complex backgrounds and translate sketches into photo-realistic images.
Figure 3: Reconstruction results for training (left) and unseen (right) identities. Specifically, we extract features of original images by using a FR model and feed them to Vec2Face for reconstruction. We observe that the reconstructed images still maintain the same identity while removing some image borders and backgrounds and transferring sketches into photo-realistic images.
3.2 Inference: Sampling vectors to generate identities and their face images
For Vec2Face, dissimilar input vectors will likely lead to images of different identities, allowing us to generate different identity images; controlling the perturbed vectors in a proper similarity range with their identity vector can generate face images with different attributes while preserving the identity. Thus, controlling the sampled vectors is important for dataset generation.
Sampling well-separated identity vectors. Suppose we want to generate $n$ identity images from $n$ identity vectors $v_{1},v_{2},...,v_{n}$, where $v_{i}\in\mathbb{R}^{d},i=1,2,...,n$. We ask that the similarity between any pair of identity vectors be less than a threshold $\tau$: $sim(v_{i},v_{j})\leq\tau,\hskip 5.69054pt\forall\;i\neq j$. This ensures that the generated images are of different identities. Following (Papantoniou et al., 2024), PCA is used to sample vectors. Specifically, we first compute the mean and covariance of the PCA-transformed face features, then use these statistics to sample vectors in the PCA space via multivariate normal distribution, and finally inverse transform these vectors back to the original feature space to obtain new identity vectors. Newly obtained identity vectors that have similarity $\leq\tau=0.3$ with all existing vectors are kept, otherwise they are dropped. Interestingly, we notice that, in the high-dimensional (512 in this paper) feature space, most of the randomly sampled vectors fit the condition due to sparsity. In fact, only 1.7% of the sampled 300k vectors are filtered out.
Perturbing identity vectors for image creation. After obtaining $n$ identity vectors $v_{1},v_{2},...,v_{n}$ , we sample $m$ vectors for each identity vector by perturbation: